# Latent Factor Analysis in Short Panels

### v2

@@ E.7 Proofs of Lemmas 4-10
Proof of Lemma 9: By the root-$n$ consistency of the FA estimators, $\hat{z}^{*}_{m,n}=z^{*}_{m,n}+O_{p}(\frac{b_{m,n}}{\sqrt{n}})$, uniformly in $m$, where $z^{*}_{m,n}=\sum_{i\in I_{m}}G^{\prime}V_{\varepsilon}^{-1}\varepsilon_{i}\varepsilon_{i}^{\prime}V_{\varepsilon}^{-1}G=\sum_{i\in I_{m}}Q^{\prime}e_{i}e^{\prime}_{i}Q$. Under the condition $\frac{1}{n^{3/2}}\sum_{m=1}^{J_{n}}b_{m,n}^{2}=\sqrt{n}\sum_{m=1}^{J_{n}}B_{m,n}^{2}=o(1)$, we have $\hat{\Omega}_{Z^{*}}=\frac{1}{n}\sum_{m=1}^{J_{n}}E[vech(z^{*}_{m,n})vech(z^{*}_{m,n})^{\prime}]+o_{p}(1)$, up to pre- and post-multiplication by an orthogonal matrix. Moreover, $vech(z^{*}_{m,n})=\boldsymbol{R}^{\prime}[\sum_{i\in I_{m}}vech(e_{i}e_{i}^{\prime})]=\frac{1}{2}\boldsymbol{R}^{\prime}A_{T}^{\prime}[\sum_{i\in I_{m}}(e_{i}\otimes e_{i})]$, and $\sum_{i\in I_{m}}(e_{i}\otimes e_{i})=\sum_{a,b}\sigma_{a,b}(w_{a}\otimes w_{b})$. Thus, we get $E\left[vech(z^{*}_{m,n})vech(z^{*}_{m,n})^{\prime}\right]=\frac{1}{4}\boldsymbol{R}^{\prime}A_{T}^{\prime}\left\{\sum_{a,b,c,d\in I_{m}}\sigma_{a,b}\sigma_{c,d}E[(w_{a}\otimes w_{b})(w_{c}\otimes w_{d})^{\prime}]\right\}A_{T}\boldsymbol{R}$. The non-zero contributions to the term in the curly brackets come from the combinations with $a=b=c=d$, $a=b\neq c=d$, $a=c\neq b=d$ and $a=d\neq b=c$, yielding: $\sum_{a,b,c,d}\sigma_{a,b}\sigma_{c,d}E[(w_{a}\otimes w_{b})(w_{c}\otimes w_{d})^{\prime}]=$ $\sum_{a}\sigma_{a,a}^{2}E[(w_{a}w_{a}^{\prime})\otimes(w_{a}w_{a}^{\prime})]+(\sum_{a\neq c}\sigma_{a,a}\sigma_{c,c})vec(I_{T})vec(I_{T})^{\prime}+(\sum_{a\neq b}\sigma_{a,b}^{2})(I_{T^{2}}+K_{T,T})=\sum_{a}[\sigma_{a,a}^{2}V(w_{a}\otimes w_{a})]+(\sum_{a}\sigma_{a,a})^{2}vec(I_{T})vec(I_{T})^{\prime}+(\sum_{a\neq b}\sigma_{a,b}^{2})(I_{T^{2}}+K_{T,T})$. Then, using $w_{a}\otimes w_{a}=A_{T}vech(w_{a}w_{a}^{\prime})$, we get $\frac{1}{4}A_{T}^{\prime}\left\{\sum_{a,b,c,d\in I_{m}}\sigma_{a,b}\sigma_{c,d}E[(w_{a}\otimes w_{b})(w_{c}\otimes w_{d})^{\prime}]\right\}A_{T}=\sum_{a}[\sigma_{a,a}^{2}V(vech(w_{a}w_{a}^{\prime}))]+(\sum_{a}\sigma_{a,a})^{2}vech(I_{T})vech(I_{T})^{\prime}+(\sum_{a\neq b}\sigma_{a,b}^{2})I_{\frac{T(T+1)}{2}}$. Then, since $\frac{1}{n}\sum_{i=1}^{n}\sigma_{i,i}^{2}V[vech(w_{i}w_{i}^{\prime})]=D_{n}$, where matrix $D_{n}$ is defined in Assumption A.6, we get $\hat{\Omega}_{Z^{*}}=\boldsymbol{R}^{\prime}\tilde{\Xi}_{n}\boldsymbol{R}+o_{p}(1)$, where $\tilde{\Xi}_{n}=D_{n}+(q_{n}+\xi_{n})vech(I_{T})vech(I_{T})^{\prime}+\kappa_{n}I_{\frac{T(T+1)}{2}}$. Moreover, under Assumption 3, and singling out parameter $q_{n}$ along the diagonal, we have $D_{n}=[\psi_{n}(0)-2q_{n}]D(0)+\sum_{h=1}^{T-1}\psi_{n}(h)[\tilde{D}(h)+\bar{D}(h)]+q_{n}I_{T(T+1)/2}$. The conclusion follows.
@@
Proof of Lemma 10: We use $vec(S)=A_{m}vech(S)$, where the $m^{2}\times\frac{1}{2}m(m+1)$ matrix $A_{m}$ is such that: (i) $A_{m}^{\prime}A_{m}=2I_{\frac{1}{2}m(m+1)}$, (ii) $K_{m,m}A_{m}=A_{m}$, where $K_{m,m}$ is the commutation matrix for order $m$, and (iii) $A_{m}A_{m}^{\prime}=I_{m^{2}}+K_{m,m}$ (see the proof of Proposition 2 and also Theorem 12 in Magnus, Neudecker (2007) Chapter 2.8). Then, $vech(S)=\frac{1}{2}A_{m}^{\prime}vec(S)$ by property (i), and $vech(O^{-1}SO)=\frac{1}{2}A_{m}^{\prime}vec(O^{-1}SO)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})vec(S)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})A_{m}vech(S),$ for all symmetric matrix $S$. It follows $\mathscr{R}(O)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})A_{m}$. Moreover, by properties (i)-(iii), we have (a) $\mathscr{R}(I_{m})=I_{\frac{1}{2}m(m+1)}$, (b) $\mathscr{R}(O_{1})\mathscr{R}(O_{2})=\frac{1}{4}A_{m}^{\prime}(O_{1}^{\prime}\otimes O_{1}^{\prime})A_{m}A_{m}^{\prime}(O_{2}^{\prime}\otimes O_{2}^{\prime})A_{m}=\frac{1}{4}A_{m}^{\prime}(O_{1}^{\prime}\otimes O_{1}^{\prime})(I_{m^{2}}+K_{m,m})(O_{2}^{\prime}\otimes O_{2}^{\prime})A_{m}=\frac{1}{4}A_{m}^{\prime}(O_{1}^{\prime}O_{2}^{\prime}\otimes O_{1}^{\prime}O_{2}^{\prime})(I_{m^{2}}+K_{m,m})A_{m}=\frac{1}{2}A_{m}^{\prime}[(O_{2}O_{1})^{\prime}\otimes(O_{2}O_{1})^{\prime}]A_{m}=\mathscr{R}(O_{2}O_{1}),$ and thus (c) $[\mathscr{R}(O)]^{-1}=\mathscr{R}(O^{-1})$.
F Numerical checks of conditions (7) of Proposition 5
[…]
F.1 Calibration of $\bar{\nu}$, $\underline{\lambda}$ and $\bar{\lambda}$
@@
To calibrate the bounds $\bar{\nu}$, $\underline{\lambda}$ and $\bar{\lambda}$ with realistic values, we run the following numerical experiment. For $T=20$ and $k=7$, we simulate $10,000$ draws from random $T\times k$ matrix $\tilde{F}$ such that $vec(\tilde{F})\sim N(0,I_{Tk})$ and set $F=V_{\varepsilon}^{1/2}U\Gamma^{1/2}$, $U=\tilde{F}(\tilde{F}^{\prime}\tilde{F})^{-1/2}$, $G=V_{\varepsilon}^{1/2}Q$, $Q=\tilde{Q}(\tilde{Q}^{\prime}\tilde{Q})^{-1/2}$, $\tilde{Q}$ are the first $T-k$ columns of $I_{T}-UU^{\prime}$, for $V_{\varepsilon}=diag(V_{\varepsilon,11},...,V_{\varepsilon,TT})$, with $V_{\varepsilon,tt}=1.5$ for $t=1,...,10$, and $V_{\varepsilon,tt}=0.5$ for $t=11,...,20$, and $\Gamma=Tdiag(4,3.5,3,2.5,2,1.5,1)$, $c_{k+1}=10T$, and $\xi_{k+1}=e_{1}$. With these choices, the “signal-to-noise" $\frac{1}{T}F_{j}^{\prime}V_{\varepsilon}^{-1}F_{j}$ for the seven factors $j=1,...,7$ are $4,3.5,3,2.5,2,2.5,1$, and the “signal-to-noise" for the weak factor is $\frac{1}{T}F_{k+1}^{\prime}V_{\varepsilon}^{-1}F_{k+1}=10n^{-1/2}$. Moreover, the errors follow the ARCH model of Section E.5.3 (i) with ARCH parameters either (a) $\alpha_{i}=0.2$ for all $i$, or (b) $\alpha_{i}=0.5$ for all $i$, and $q=4$, and $\kappa=0$ (cross-sectional independence). The choices $\alpha_{i}=0.2$, $0.5$ both meet the condition $13\alpha_{i}^{2}<$ 1 ensuring the existence of fourth-order moments. Moreover, with $q-1=3$, we have a cross-sectional variance of the $\sigma_{ii}$ that is three times larger than the mean (normalized to $1$). For each draw, we compute the $df=71$ non-zero eigenvalues and associated eigenvectors of $\Omega_{\bar{Z}^{*}}$, and the values of parameters $\nu_{j}$ and $\lambda_{j}$. In our simulations (a) with $\alpha_{i}=0.2$, the draws of $\max_{j=1,...,df}\nu_{j}$ range between $0.21$ and $0.30$, with $95\%$ quantile equal to $0.28$, while the $5\%$ and $95\%$ quantiles of the $\lambda_{j}$ are $0.13$ and $7.65$. Instead, (b) with $\alpha_{i}=0.5$, the $\max_{j=1,...,df}\nu_{j}$ range between $0.70$ and $0.79$, with $95\%$ quantile equal to $0.77$, and the $5\%$ and $95\%$ quantiles of the $\lambda_{j}$ are $0.12$ and $6.64$. To get further insights in the choice of parameters $\bar{\nu},\underline{\lambda},\bar{\lambda}$, we also consider the values implied by the FA estimates in our empirical analysis. Here, when testing for the last retained $k$ in a given subperiod, the median across subperiods of $\max_{j=1,...,df}\nu_{j}$ is $0.76$, and smaller than about $0.90$ in most subperiods. Similarly, assuming $c_{k+1}=10T$ and $\xi_{k+1}=e_{1}$ as above, the median values of the smallest and the largest estimated $\lambda_{j}$ are $0.0024$ and $5.84$. Inspired by these findings, we set $\bar{\lambda}=7$, and consider $\bar{\nu}=0.2$, $0.7$, $0.9$, $0.99$, and $\underline{\lambda}=0$, $0.1$, $0.5$, $1$, to get realistic settings with different degrees of dissimilarity from the case with serially uncorrelated squared errors (increasing with $\bar{\nu}$), and separation of the alternative hypothesis from the null hypothesis (increasing with $\underline{\lambda}$).
F.2 Results with Monte Carlo draws
[…]
$df$ 2 3 4 5 6 7 8 9 10 11 12

### v3

@@ B.7 Proofs of Lemmas 5-11
Proof of Lemma 10: By the root-$n$ consistency of the FA estimators, $\hat{z}^{*}_{m,n}=z^{*}_{m,n}+O_{p}(\frac{b_{m,n}}{\sqrt{n}})$, uniformly in $m$, where $z^{*}_{m,n}=\sum_{i\in I_{m}}G^{\prime}V_{\varepsilon}^{-1}\varepsilon_{i}\varepsilon_{i}^{\prime}V_{\varepsilon}^{-1}G=\sum_{i\in I_{m}}Q^{\prime}e_{i}e^{\prime}_{i}Q$. Under the condition $\frac{1}{n^{3/2}}\sum_{m=1}^{J_{n}}b_{m,n}^{2}=\sqrt{n}\sum_{m=1}^{J_{n}}B_{m,n}^{2}=o(1)$, we have $\hat{\Omega}_{Z^{*}}=\frac{1}{n}\sum_{m=1}^{J_{n}}E[vech(z^{*}_{m,n})vech(z^{*}_{m,n})^{\prime}]+o_{p}(1)$, up to pre- and post-multiplication by an orthogonal matrix. Moreover, $vech(z^{*}_{m,n})=\boldsymbol{R}^{\prime}[\sum_{i\in I_{m}}vech(e_{i}e_{i}^{\prime})]=\frac{1}{2}\boldsymbol{R}^{\prime}A_{T}^{\prime}[\sum_{i\in I_{m}}(e_{i}\otimes e_{i})]$, and $\sum_{i\in I_{m}}(e_{i}\otimes e_{i})=\sum_{a,b}\sigma_{a,b}(w_{a}\otimes w_{b})$. Thus, we get $E\left[vech(z^{*}_{m,n})vech(z^{*}_{m,n})^{\prime}\right]=\frac{1}{4}\boldsymbol{R}^{\prime}A_{T}^{\prime}\left\{\sum_{a,b,c,d\in I_{m}}\sigma_{a,b}\sigma_{c,d}E[(w_{a}\otimes w_{b})(w_{c}\otimes w_{d})^{\prime}]\right\}A_{T}\boldsymbol{R}$. The non-zero contributions to the term in the curly brackets come from the combinations with $a=b=c=d$, $a=b\neq c=d$, $a=c\neq b=d$ and $a=d\neq b=c$, yielding: $\sum_{a,b,c,d}\sigma_{a,b}\sigma_{c,d}E[(w_{a}\otimes w_{b})(w_{c}\otimes w_{d})^{\prime}]=$ $\sum_{a}\sigma_{a,a}^{2}E[(w_{a}w_{a}^{\prime})\otimes(w_{a}w_{a}^{\prime})]+(\sum_{a\neq c}\sigma_{a,a}\sigma_{c,c})vec(I_{T})vec(I_{T})^{\prime}+(\sum_{a\neq b}\sigma_{a,b}^{2})(I_{T^{2}}+K_{T,T})=\sum_{a}[\sigma_{a,a}^{2}V(w_{a}\otimes w_{a})]+(\sum_{a}\sigma_{a,a})^{2}vec(I_{T})vec(I_{T})^{\prime}+(\sum_{a\neq b}\sigma_{a,b}^{2})(I_{T^{2}}+K_{T,T})$. Then, using $w_{a}\otimes w_{a}=A_{T}vech(w_{a}w_{a}^{\prime})$, we get $\frac{1}{4}A_{T}^{\prime}\left\{\sum_{a,b,c,d\in I_{m}}\sigma_{a,b}\sigma_{c,d}E[(w_{a}\otimes w_{b})(w_{c}\otimes w_{d})^{\prime}]\right\}A_{T}=\sum_{a}[\sigma_{a,a}^{2}V(vech(w_{a}w_{a}^{\prime}))]+(\sum_{a}\sigma_{a,a})^{2}vech(I_{T})vech(I_{T})^{\prime}+(\sum_{a\neq b}\sigma_{a,b}^{2})I_{\frac{T(T+1)}{2}}$. Then, since $\frac{1}{n}\sum_{i=1}^{n}\sigma_{i,i}^{2}V[vech(w_{i}w_{i}^{\prime})]=D_{n}$, where matrix $D_{n}$ is defined in Assumption 8 we get $\hat{\Omega}_{Z^{*}}=\boldsymbol{R}^{\prime}\tilde{\Xi}_{n}\boldsymbol{R}+o_{p}(1)$, where $\tilde{\Xi}_{n}=D_{n}+(q_{n}+\xi_{n})vech(I_{T})vech(I_{T})^{\prime}+\kappa_{n}I_{\frac{T(T+1)}{2}}$. Moreover, under Assumption 11, and singling out parameter $q_{n}$ along the diagonal, we have $D_{n}=[\psi_{n}(0)-2q_{n}]D(0)+\sum_{h=1}^{T-1}\psi_{n}(h)[\tilde{D}(h)+\bar{D}(h)]+q_{n}I_{T(T+1)/2}$. The conclusion follows.
@@
Proof of Lemma 11: We use $vec(S)=A_{m}vech(S)$, where the $m^{2}\times\frac{1}{2}m(m+1)$ matrix $A_{m}$ is such that: (i) $A_{m}^{\prime}A_{m}=2I_{\frac{1}{2}m(m+1)}$, (ii) $K_{m,m}A_{m}=A_{m}$, where $K_{m,m}$ is the commutation matrix for order $m$, and (iii) $A_{m}A_{m}^{\prime}=I_{m^{2}}+K_{m,m}$ (see the proof of Proposition 1 and also Theorem 12 in Magnus, Neudecker (2007) Chapter 2.8). Then, $vech(S)=\frac{1}{2}A_{m}^{\prime}vec(S)$ by property (i), and $vech(O^{-1}SO)=\frac{1}{2}A_{m}^{\prime}vec(O^{-1}SO)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})vec(S)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})A_{m}vech(S),$ for all symmetric matrix $S$. It follows $\mathscr{R}(O)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})A_{m}$. Moreover, by properties (i)-(iii), we have (a) $\mathscr{R}(I_{m})=I_{\frac{1}{2}m(m+1)}$, (b) $\mathscr{R}(O_{1})\mathscr{R}(O_{2})=\frac{1}{4}A_{m}^{\prime}(O_{1}^{\prime}\otimes O_{1}^{\prime})A_{m}A_{m}^{\prime}(O_{2}^{\prime}\otimes O_{2}^{\prime})A_{m}=\frac{1}{4}A_{m}^{\prime}(O_{1}^{\prime}\otimes O_{1}^{\prime})(I_{m^{2}}+K_{m,m})(O_{2}^{\prime}\otimes O_{2}^{\prime})A_{m}=\frac{1}{4}A_{m}^{\prime}(O_{1}^{\prime}O_{2}^{\prime}\otimes O_{1}^{\prime}O_{2}^{\prime})(I_{m^{2}}+K_{m,m})A_{m}=\frac{1}{2}A_{m}^{\prime}[(O_{2}O_{1})^{\prime}\otimes(O_{2}O_{1})^{\prime}]A_{m}=\mathscr{R}(O_{2}O_{1}),$ and thus (c) $\mathscr{R}(O)]^{-1}=\mathscr{R}(O^{-1})$.
C Numerical checks of conditions (5) of Proposition 4
In this section, we check numerically the validity of Inequalities (5) for given $df$, $\lambda_{j}$, $\nu_{j}$, and m=3,.,Mm=3,....,M, for a large bound $M$. The idea is to compute the frequency of the LHS of (5) becoming strictly negative over a large number of potential values of $\lambda_{j}$ and $\nu_{j}$, j=1,,dfj=1,...,df, for any given df>1df>1.181818From Footnote 2, we know that Inequalities (5) are automatically met with $df=1$. A given value of $df$ may result from several different combinations of $T$ and $k$, while a given $T$ implies different values of $df$ depending on $k$. For instance, $df=2$ applies with $(T,k)=(4,1)$, $(8,4)$, and $(13,8)$, among other combinations. For $T=20$, the tests for k=1,2,,14k=1,2,...,14 yield df=170,151,133,116,100,85,71,58,46,35,25,16,8,1df=170,151,133,116,100,85,71,58,46,35,25,16,8,1, respectively. Table 2 provides those frequencies for m=3,,16m=3,...,16 (cumulatively), with $\lambda_{j}$ uniformly drawn in $\underline{\lambda},\bar{\lambda}]$ for j=1,,dfj=1,...,df, and with $\nu_{1}=0$  [^19] This normalization results from ranking the eigenvalues $\mu_{j}$, so that $\mu_{1}$ is the smallest one. and $\nu_{j}$ uniformly drawn in $0,\bar{\nu}]$, for j=2,,dfj=2,...,df, and different combinations of bounds $\underline{\lambda}$, $\bar{\lambda}$, $\bar{\nu}$, and degrees of freedom df=2,,12df=2,...,12. Each frequency is computed from $10^{8}$ draws of $\lambda_{j}$ and $\nu_{j}$, j=1,,dfj=1,...,df. In the SMC, we also report a table of frequencies for large grids of equally-spaced points in $\underline{\lambda},\bar{\lambda}]^{df}\times[0,\bar{\nu}]^{df-1}$, which corroborate the findings of this section.
C.1 Calibration of $\bar{\nu}$, $\underline{\lambda}$ and $\bar{\lambda}$
@@
To calibrate the bounds $\bar{\nu}$, $\underline{\lambda}$ and $\bar{\lambda}$ with realistic values, we run the following numerical experiment. For $T=20$ and $k=7$, we simulate $10,000$ draws from random $T\times k$ matrix $\tilde{F}$ such that $vec(\tilde{F})\sim N(0,I_{Tk})$ and set $F=V_{\varepsilon}^{1/2}U\Gamma^{1/2}$, $U=\tilde{F}(\tilde{F}^{\prime}\tilde{F})^{-1/2}$, $G=V_{\varepsilon}^{1/2}Q$, $\tilde{Q}(\tilde{Q}^{\prime}\tilde{Q})^{-1/2}, Q~\tilde{Q}$ are the first TkT-k columns of $I_{T}-UU^{\prime}$, for $V_{\varepsilon}=diag(V_{\varepsilon,11},...,V_{\varepsilon,TT})$, with $V_{\varepsilon,tt}=1.5$ for t=1,,10t=1,...,10, and $V_{\varepsilon,tt}=0.5$ for t=11,,20t=11,...,20, and $\Gamma=Tdiag(4,3.5,3,2.5,2,1.5,1)$, $c_{k+1}=10T$, and $\xi_{k+1}=e_{1}$. With these choices, the “signal-to-noise" $\frac{1}{T}F_{j}^{\prime}V_{\varepsilon}^{-1}F_{j}$ for the seven factors j=1,,7j=1,...,7 are $4,3.5,3,2.5,2,2.5,1$, and the “signal-to-noise" for the weak factor is $\frac{1}{T}F_{k+1}^{\prime}V_{\varepsilon}^{-1}F_{k+1}=10n^{-1/2}$. Moreover, the errors follow the ARCH model of Section B.5.3 (i) with ARCH parameters either (a) $\alpha_{i}=0.2$ for all $i$, or (b) $\alpha_{i}=0.5$ for all $i$, and $q=4$, and $\kappa=0$ (cross-sectional independence). The choices $\alpha_{i}=0.2$, $0.5$ both meet the condition $3\alpha_{i}^{2}<1$ ensuring the existence of fourth-order moments. Moreover, with q1=3q-1=3, we have a cross-sectional variance of the $\sigma_{ii}$ that is three times larger than the mean (normalized to 11). For each draw, we compute the $df=71$ non-zero eigenvalues and associated eigenvectors of $\Omega_{\bar{Z}^{*}}$, and the values of parameters $\nu_{j}$ and $\lambda_{j}$. In our simulations (a) with $\alpha_{i}=0.2$, the draws of $\max_{j=1,...,df}\nu_{j}$ range between $0.21$ and $0.30$, with $95\%$ quantile equal to $0.28$, while the $5\%$ and $95\%$ quantiles of the $\lambda_{j}$ are $0.13$ and $7.65$. Instead, (b) with $\alpha_{i}=0.5$, the $\max_{j=1,...,df}\nu_{j}$ range between $0.70$ and $0.79$, with $95\%$ quantile equal to $0.77$, and the $5\%$ and $95\%$ quantiles of the $\lambda_{j}$ are $0.12$ and $6.64$. To get further insights in the choice of parameters $\bar{\nu},\underline{\lambda},\bar{\lambda}$, we also consider the values implied by the FA estimates in our empirical analysis. Here, when testing for the last retained $k$ in a given subperiod, the median across subperiods of $\max_{j=1,...,df}\nu_{j}$ is $0.76$, and smaller than about $0.90$ in most subperiods. Similarly, assuming $c_{k+1}=10T$ and $\xi_{k+1}=e_{1}$ as above, the median values of the smallest and the largest estimated $\lambda_{j}$ are $0.0024$ and $5.84$. Inspired by these findings, we set $\bar{\lambda}=7$, and consider $\bar{\nu}=0.2$, $0.7$, $0.9$, $0.99$, and $\underline{\lambda}=0$, $0.1$, $0.5$, 11, to get realistic settings with different degrees of dissimilarity from the case with serially uncorrelated squared errors (increasing with $\bar{\nu}$), and separation of the alternative hypothesis from the null hypothesis (increasing with $\underline{\lambda}$).
C.2 Results with Monte Carlo draws
In Table 2, the entries are nil for $\bar{\nu}$ sufficiently small and $\underline{\lambda}$ sufficiently large, suggesting that the AUMPI property holds for those cases that are closer to the setting with uncorrelated squared errors and sufficiently separated from the null hypothesis. Violations of Inequalities (5) concern df=2,3,4,5df=2,3,4,5.202020A given number of simulated draws become increasingly sparse when considering larger values of $df$, which makes the exploration of the parameter space more challenging in those cases. However, unreported theoretical considerations show via an asymptotic approximation that the monotone likelihood property holds for $df\rightarrow\infty$ since the limiting distribution is then Gaussian. This finding resonates with the absence of violations in Table 2 for the larger values of $df$. Let us focus on the setting with $\bar{\nu}=0.7$ and $\underline{\lambda}=0.1$. We find $3752$ violations of Inequalities (5) out of $10^{8}$ simulations, all occurring for $df=2$, except $65$ for $df=3$. For those draws violating Inequalities (5) for $df=2$, a closer inspection shows that (a) they feature values $\nu_{2}$ close to upper bound $\bar{\nu}=0.7$, and values of $\lambda_{2}$ close to lower bound $\underline{\lambda}=0.1$, and (b) several of them yield non-monotone density ratios $\frac{f(z;\lambda_{1},\lambda_{2})}{f(z;0,0)}$, with the non-monotonicity region corresponding to large values of $z$. As an illustration, let us take the density ratio for $df=2$ with $\nu_{2}=0.666$, $\lambda_{1}=1.372$, and $\lambda_{2}=0.130$. Here, the eigenvalues of the covariance matrix are $\mu_{1}=1$ (by normalization) and $\mu_{2}=(1-\nu_{2})^{-1}=2.994$, and the non-centrality parameter $\lambda_{2}$ is small. The quantiles of the asymptotic distribution under the null hypothesis for asymptotic size $\alpha=20\%,10\%,5\%,1\%,0.1\%$ are $9.3,12.8,16.2,24.5,36.5$. Non-monotonicity applies for $z\geq 16$. The optimal rejection regions $\{\frac{f(z;\lambda_{1},\lambda_{2})}{f(z;0,0)}\geq C\}$ correspond to those of the LR test $\{z\geq\tilde{C}\}$, e.g., for asymptotic levels such as $\alpha=20\%$, but not for $\alpha=5\%$ or smaller. Indeed, in the latter cases, because of non-monotonicity of the density ratio, the optimal rejection regions are finite intervals in argument $z$. With $\bar{\nu}=0.7$, we do not find violations with $\underline{\lambda}=0.5$ or larger.
dfdf 2 3 4 5 6 7 8 9 10 11 12